Define the overall heat transfer coefficient? Derive the relation among overall heat transfer coefficient and individual heat transfer coefficients?


1
Expert's answer
2021-09-03T07:21:01-0400

The overall heat transfer

The overall heat transfer refers to how well heat is conducted over a series of resilient mediums. Its units are the W/(m2°C) [Btu/(hr-ft2°F)].

Overall heat transfer coefficient) is given by, (1/UA) = (1/Uo*Ao) = (1/Ui*Ai) = { (1 / (hr * Ar)) + (Rf,r / Ar) + ( ( ln(Do/Di)) / (2 * PI * Kwall * Ltube)) + (Rf,liq / Aliq) + ( 1/ (hliq * Aliq))}

Steps: 

1.     Start your calculation by approximately selecting sizes of tubes involved

2.     Check in the final stage, whether they meet the criteria. If not again calculate with a different selection of sizes involved.

Things needed: 

1. Required cooling or heating capacity 

2. Overall heat transfer coefficient of heat exchanger based on thermal and fluids properties plus materials of heat exchangers involved. 

3. Log mean temperature difference of both fluids.